# Structural changes in the Lennard-Jones supercooled liquid and ideal glass: an improved integral equation for the replica method

### v1

@@ I Introduction
Recent state-of-the-art computer studies and theoretical analysis have provided support to the hypothesis that supercooled liquids, carefully maintained in equilibrium conditions at low temperatures, can undergo a discontinuous liquid-glass phase transition, questioning in this way the assumption that the glass formation is essentially a dynamic process accompanied by small structural changes Guiselin2022 . In this context, a fundamentally thermodynamic perspective of the glass transition was provided by the so-called “random first-order transition” (RFOT) theory Wolynes , that includes the elegant “replica method”, which purpose is to locate the ideal-glass transition in the density-temperature plane Franz1995 ; Mezard1996 ; Franz1997 ; Franz1998 ; Cardenas1998 ; Cardenas1999 ; Coluzzi1999 . Such a static approach involves Integral Equation Theories (IET) BomontACP2008 ; Hansen2013 ; Janssen2024 , and gave rise, since then, to many studies (see Ref. 30 for an extensive review). The replica method requires the use of an appropriate order parameter $Q$, which quantifies the overlap between two coupled clones of a system and therefore can distinguish between the equilibrium liquid and the glass phases. The overlap is also useful above the glass transition, since it has been suggested that the RFOT might be preceded by a “precursor transition” occurring at higher temperatures, signaling the coexistence of two phases corresponding to low and high values of $Q$. The low-$Q$ phase is associated with the supercooled liquid, while the high-$Q$ phase is for the glass. Later on, a simplified but equivalent scheme was proposed by two of us in collaboration with Jean-Pierre Hansen BomontEPL2014 and, since then, successfully applied to a variety of interaction models BomontJCP2014 ; BomontMOLPHYS2015 ; BomontPRE2015 ; Bomont2017 ; BomontJCP2017 ; BomontJCP2019 ; BomontPRE2022 ; BomontJCP2024 . In particular, we compared the predictions coming from our simplified scheme BomontEPL2014 with those obtained within the original theory Franz2013 for two different interaction models, namely soft spheres BomontJCP2019 and Lennard-Jones BomontJCP2024 . Thereby, we showed that, under strictly identical conditions, the two approaches provide identical results for both models, as for the dynamical transition.
@@
In the present work, we assume that thermal equilibrium can be maintained at an arbitrary low temperature, and we study how deeply a liquid can be supercooled within the framework of IET. In particular, we aim to investigate the fate of structural correlations within the deeply supercooled liquid and the ideal-glass phase. To our purposes, we study the Lennard-Jones model by a replicated version of the thermodynamically self-consistent hybridized-mean-spherical approximation theory (R-HMSA) Zerah , combined with the Optimized Division Scheme for the pair interaction proposed by two of us Bomont2001 , that is known to provide accurate predictions for the structure and thermodynamics of simple liquids Bomont1997 ; Bomont1998 . At a first stage, upon cooling the system along several isochores, thermodynamically self-consistent results are found in ranges of low temperatures never reached before by any other computational tool, either in the liquid or in the ideal-glass phases. Within this regime, we identify the upper boundary, given by the dynamical transition temperature $T_{\rm D}$, for the existence of the ideal-glass phase. Then, by free energy calculations, we locate the critical temperature $T_{\rm cr}<T_{\rm D}$ below which the ideal-glass phase is the stable one. Finally, this offers us the opportunity to identify in both phases a density-dependent low temperature threshold $T_{0}<T_{\rm cr}$ across which a clear-cut structural change occurs. Our choice for the pure Lennard-Jones, in comparison for example with a standard glass former such as the Kob-Andersen mixture Kob1994 ; Kob1995A ; Kob1995B is dictated not only by the possibility to characterize the glass behavior of perhaps the most exhaustively studied model for the fluid state. In fact, since in our formalism the study of a pure fluid involves the introduction of a symmetric binary mixture, starting from a binary mixture at the outset, would drastically increase the complexity of the scheme, requiring in particular several different relationships to let us predict the pair structure of the fluid in a thermodynamically consistent manner.
@@ II.2 HMSA scheme and ODS splitting procedure
For a given state point, the equilibrium pair structure of the symmetric binary mixture is characterized by two pair distribution functions $g_{11}(r)\equiv g_{22}(r)$ and $g_{12}(r)$ which can be calculated approximately by solving the two coupled integral equations resulting from the two Ornstein-Zernike relations between the $g_{ij}(r)$ and the direct and indirect correlation functions $c_{ij}(r)$, $\gamma_{ij}(r)=g_{ij}(r)-1-c_{ij}(r)$ respectively, upon use of approximate closure relations Hansen2013 . To our purposes, we choose the HMSA closure Zerah , whose radial distribution function for the symmetric binary-mixture reads
@@ III Interplay between supercooled liquid and ideal glass
Predictions for $f_{\rm ex}^{\ast\rm L}$ and $f_{\rm ex}^{\ast\rm G}$, with generically $f_{\rm ex}^{\ast}=F_{\rm ex}^{\ast}/N\varepsilon$ — as deduced accordingly to the thermodynamic integration schemes described Sect. II.3 — are reported along two different isochores in Fig. 4. As seen, the free energies corresponding to the liquid and ideal-glass phases are very close to each other, as also observed in a previous HNC study BomontPRE2015 . Along each isochore, $f_{\rm ex}^{\ast\rm L}$ lies below $f_{\rm ex}^{\ast\rm G}$, until a crossing is clearly observed at a temperature $T^{\ast}_{\mathrm{cr}}$, pointing to a weak first-order transition to the ideal glass, with the latter being stable for $T^{\ast}<T^{\ast}_{\mathrm{cr}}$. Our predictions for $T^{\ast}_{\mathrm{cr}}$ are also shown in Fig. 3 (green pluses) along with $T^{\ast}_{\rm D}$. Numerical values are listed in Table 1.
@@
Finally, we mention  — among various attempts to study theoretically the high-density/low-temperature behavior of LJ —  an improved version of the Modified HNC theory Rosenfeld , developed by our late colleague and friend Mimmo Gazzillo Gazzillo1993 . In that work, the authors carried out a positive assessment of their structural predictions against simulation data. However, even in that case, the theoretical approach failed to attain a very low temperature regime. For instance, the lowest temperature reached at $\rho^{\ast}=0.95$ in Ref. 71 was $T^{\ast}_{\rm min}=0.233$, to be compared with the present $T^{\ast}_{\rm min}(\rho^{\ast}=0.95)=0.007$.
As far as simulation studies are concerned, it is well recognized that the investigation of glass-forming systems suffer from the rapidly growing relaxation times near the glass transition, which historically have limited investigations into the regime of very mild supercooling Ediger1996 . In this context, we recall that the very first attempt to investigate by molecular dynamics the glass state of the LJ fluid, dates back to almost fifty years ago: in Ref. 73, Rahman and co-workers were in fact able to generate an amorphous sample by means of a rapid, isochoric quenching from the liquid phase. Since then, a variety of methods — as for instance simulated tempering and cluster Monte Carlo Dress1995  — were introduced to deal with issues related to slow equilibration. More recently, swap Monte Carlo schemes — in which the standard Metropolis sampling is “augmented” by swap moves between physically distant particles of different species — were shown to substantially speed-up the equilibration in certain glass-forming mixtures, allowing for the study of glassy properties previously out of reach swap . Nevertheless, these huge computational efforts are devoted to the study of polydisperse models only, and even in this case, the technique is thought to be “too a primitive approach” Berthier2023 requiring substantial improvements to firmly test the validity of the RFOT theory.
@@ IV Structural properties in the low-temperature regime
Figure 5: Comparison between $g(r)$ in the liquid (blue) and $g_{11}(r)$ in the ideal glass (red) phases for $\rho^{\ast}=0.9$ at $T^{\ast}=0.2$ (full lines) and $T^{\ast}=0.1$ (dashed lines, shifted by one along the vertical axis, for clarity)
@@
Figure 6: Variation of $g(r)$ (A, B) and $g_{11}(r)$ (C, D) across $T_{\mathrm{0}}^{\ast}$ at $\rho^{\ast}=0.90$ (A, C) and $\rho^{\ast}=0.95$ (B, D); temperatures in the legend. The dashed line in A refers to HMSA predictions coupled with WCA, at $T^{\ast}=0.2$. Circles in B refer to MD data at $T^{\ast}=0.233$ of Ref. 71. For clarity, in each panel, all curves but the lowest one are shifted by two along the vertical axis.
@@
In Fig. 7 $g_{11}(r)$ and $g_{12}(r)$ for $\rho^{\ast}=0.90$ at $T^{\ast}=0.0625$ are shown. As seen, both functions oscillate in phase, according to the expectation that a unique underlying disordered equilibrium configuration of the ideal-glass exists, around which the atoms of the two replicas vibrate. Moreover, $g_{12}(r)$ signals as well the local structural change occurring within the ideal-glass, with the presence of the additional peak at about $r/\sigma=1.45$, in between the original first two.
@@
To close this discussion, HMSA/WCA $g(r)$ at $\rho^{\ast}=0.90$ at $T^{\ast}=0.2$ is shown as a dashed line in Fig. 6A. As seen, it compares well with ODS $g(r)$, but in the proximity of the first minimum. This discrepancy will progressively transform, as the temperature is lowered, into a negative “correlation hole”, eventually triggering the failure of the HMSA/WCA numerical integration algorithm. Finally, the study in Ref. 71 offers us the opportunity to positively assess our theoretical predictions for $g(r)$ against molecular dynamics results, at least in one case. This comparison is shown at $\rho^{\ast}=0.95$ in Fig. 6B, for $T^{\ast}=0.233$, i.e. the lowest temperature attained in that paper. Looking at small discrepancies in the figure, we observe that the deformation of the second peak, barely arising in the MD $g(r)$, turns out to be simply delayed to lower temperatures by HMSA. We stress once again that — to the best of our knowledge — no much more simulation structural data are available in the low-temperature regime to compare with, for the monodisperse Lennard-Jones model.

### v2

@@ I Introduction
Recent state-of-the-art computer studies and theoretical analysis have provided support to the hypothesis that supercooled liquids, carefully maintained in equilibrium conditions at low temperatures, can undergo a discontinuous liquid-glass phase transition, questioning in this way the assumption that the glass formation is essentially a dynamic process accompanied by small structural changes Guiselin2022 .
@@
In this context, a fundamentally thermodynamic perspective of the glass transition was provided by the so-called “random first-order transition” (RFOT) theory Wolynes , that includes the elegant “replica method”, which purpose is to locate the ideal-glass transition in the density-temperature plane Franz1995 ; Mezard1996 ; Franz1997 ; Franz1998 ; Cardenas1998 ; Cardenas1999 ; Coluzzi1999 . Such a static approach involves Integral Equation Theories (IET) BomontACP2008 ; Hansen2013 ; Janssen2024 , and gave rise, since then, to many studies (see Ref. 30 for an extensive review). The replica method requires the use of an appropriate order parameter $Q$, which quantifies the overlap between two coupled clones of a system and therefore can distinguish between the equilibrium liquid and the glass phases. The overlap is also useful above the glass transition, since it has been suggested that the RFOT might be preceded by a “precursor transition” occurring at higher temperatures, signaling the coexistence of two phases corresponding to low and high values of $Q$. The low-$Q$ phase is associated with the supercooled liquid, while the high-$Q$ phase is for the glass. Later on, a simplified but equivalent scheme was proposed by two of us in collaboration with Jean-Pierre Hansen BomontEPL2014 and, since then, successfully applied to a variety of interaction models BomontJCP2014 ; BomontMOLPHYS2015 ; BomontPRE2015 ; Bomont2017 ; BomontJCP2017 ; BomontJCP2019 ; BomontPRE2022 ; BomontJCP2024 .
In particular, we compared the predictions coming from our simplified scheme BomontEPL2014 with those obtained within the original theory Franz2013 for two different interaction models, namely soft spheres BomontJCP2019 and Lennard-Jones BomontJCP2024 . Thereby, we showed that, under strictly identical conditions, the two approaches provide identical results for both models, as for the dynamical transition.
@@
In the present work, we assume that thermal equilibrium can be maintained at an arbitrary low temperature, and we study how deeply a liquid can be supercooled within the framework of IET. In particular, we aim to investigate the fate of structural correlations within the deeply supercooled liquid and the ideal-glass phase. To our purposes, we study the Lennard-Jones model by a replicated version of the thermodynamically self-consistent hybridized-mean-spherical approximation theory (R-HMSA) Zerah , combined with the Optimized Division Scheme for the pair interaction proposed by two of us Bomont2001 , that is known to provide accurate predictions for the structure and thermodynamics of simple liquids Bomont1997 ; Bomont1998 . At a first stage, upon cooling the system along several isochores, thermodynamically self-consistent results are found in ranges of low temperatures never reached before by any other computational tool, either in the liquid or in the ideal-glass phases. Within this regime, we identify the upper boundary, given by the dynamical transition temperature $T_{\rm D}$, for the existence of the ideal-glass phase. Then, by free energy calculations, we locate the critical temperature $T_{\rm cr}<T_{\rm D}$ below which the ideal-glass phase is the stable one. Finally, this offers us the opportunity to identify in both phases a density-dependent low temperature threshold $T_{0}<T_{\rm cr}$ across which a clear-cut structural change occurs.
Our choice for the pure Lennard-Jones, in comparison for example with a standard glass former such as the Kob-Andersen mixture Kob1994 ; Kob1995A ; Kob1995B is dictated not only by the possibility to characterize the glass behavior of perhaps the most exhaustively studied model for the fluid state. In fact, since in our formalism the study of a pure fluid involves the introduction of a symmetric binary mixture, starting from a binary mixture at the outset, would drastically increase the complexity of the scheme, requiring in particular several different relationships to let us predict the pair structure of the fluid in a thermodynamically consistent manner.
@@ II.2 HMSA scheme and ODS splitting procedure
For a given state point, the equilibrium pair structure of the symmetric binary mixture is characterized by two pair distribution functions $g_{11}(r)\equiv g_{22}(r)$ and $g_{12}(r)$ which can be calculated approximately by solving the two coupled integral equations resulting from the two Ornstein-Zernike relations between the $g_{ij}(r)$ and the direct and indirect correlation functions $c_{ij}(r)$, $\gamma_{ij}(r)=g_{ij}(r)-1-c_{ij}(r)$ respectively, upon use of approximate closure relations Hansen2013 .
To our purposes, we choose the HMSA closure Zerah , whose radial distribution function for the symmetric binary-mixture reads
@@ III Interplay between supercooled liquid and ideal glass
Predictions for $f_{\rm ex}^{\ast\rm L}$ and $f_{\rm ex}^{\ast\rm G}$, with generically $f_{\rm ex}^{\ast}=F_{\rm ex}^{\ast}/N\varepsilon$ — as deduced accordingly to the thermodynamic integration schemes described Sect. II.3 — are reported along two different isochores in Fig. 4.
@@
As seen, the free energies corresponding to the liquid and ideal-glass phases are very close to each other, as also observed in a previous HNC study BomontPRE2015 . Along each isochore, $f_{\rm ex}^{\ast\rm L}$ lies below $f_{\rm ex}^{\ast\rm G}$, until a crossing is clearly observed at a temperature $T^{\ast}_{\mathrm{cr}}$, pointing to a weak first-order transition to the ideal glass, with the latter being stable for $T^{\ast}<T^{\ast}_{\mathrm{cr}}$. Our predictions for $T^{\ast}_{\mathrm{cr}}$ are also shown in Fig. 3 (green pluses) along with $T^{\ast}_{\rm D}$. Numerical values are listed in Table 1.
Figure 5: Nonergodicity factor $f_{\rm q}$ versus $q\sigma$ calculated by R-HNC (circles) for $\rho^{\ast}=1$ at its dynamical transition temperature $T^{\ast}=0.46$ compared to the one obtained by R-HMSA (full line) at the same state point.
Within the MCT framework, the glass transition is interpreted as a dynamical transition at a critical temperature from an ergodic (liquid) to a nonergodic phase (glass). The order parameter is the so-called nonergodicity factor $f_{\rm q}$ (the central object of MCT) defined in the reciprocal $q$-space as the long-time limit of the coherent normalized scattering function $F(q,t)$. The fact that $F(q,t\rightarrow\infty)$ does not vanish indicates that the system is in a glass phase. Whatever the potential model, it has been shown that the MCT $f_{\rm q}$ parameter is related to the counterpart of $g_{12}(r)$ in the $q$-space, namely $S_{12}(q)$, so that $f_{\rm q}=S_{12}(q)/S_{11}(q)$, where $S_{11}(q)$ is the intra-replica structure factor REVMOD2010 .
[…]
@@
Finally, we mention  — among various attempts to study theoretically the high-density/low-temperature behavior of LJ —  an improved version of the Modified HNC theory Rosenfeld , developed by our late colleague and friend Mimmo Gazzillo Gazzillo1993 . In that work, the authors carried out a positive assessment of their structural predictions against simulation data. However, even in that case, the theoretical approach failed to attain a very low temperature regime. For instance, the lowest temperature reached at $\rho^{\ast}=0.95$ in Ref. 74 was $T^{\ast}_{\rm min}=0.233$, to be compared with the present $T^{\ast}_{\rm min}(\rho^{\ast}=0.95)=0.007$.
Figure 6: Comparison between $g(r)$ in the liquid (blue) and $g_{11}(r)$ in the ideal glass (red) phases for $\rho^{\ast}=0.9$ at $T^{\ast}=0.2$ (full lines) and $T^{\ast}=0.1$ (shifted by one along the vertical axis, for clarity)
As far as simulation studies are concerned, it is well recognized that the investigation of glass-forming systems suffer from the rapidly growing relaxation times near the glass transition, which historically have limited investigations into the regime of very mild supercooling Ediger1996 . In this context, we recall that the very first attempt to investigate by molecular dynamics the glass state of the LJ fluid, dates back to almost fifty years ago: in Ref. 76, Rahman and co-workers were in fact able to generate an amorphous sample by means of a rapid, isochoric quenching from the liquid phase. Since then, a variety of methods — as for instance simulated tempering and cluster Monte Carlo Dress1995  — were introduced to deal with issues related to slow equilibration. More recently, swap Monte Carlo schemes — in which the standard Metropolis sampling is “augmented” by swap moves between physically distant particles of different species — were shown to substantially speed-up the equilibration in certain glass-forming mixtures, allowing for the study of glassy properties previously out of reach swap . Nevertheless, these huge computational efforts are devoted to the study of polydisperse models only, and even in this case, the technique is thought to be “too a primitive approach” Berthier2023 requiring substantial improvements to firmly test the validity of the RFOT theory.
@@ IV Structural properties in the low-temperature regime
Figure 7: Variation of $g(r)$ (A, B) and $g_{11}(r)$ (C, D) across $T_{\mathrm{0}}^{\ast}$ at $\rho^{\ast}=0.90$ (A, C) and $\rho^{\ast}=0.95$ (B, D); temperatures in the legend. The dashed line in A refers to HMSA predictions coupled with WCA, at $T^{\ast}=0.2$. Circles in B refer to MD data at $T^{\ast}=0.233$ of Ref. 74. For clarity, in each panel, all curves but the lowest one are shifted by two along the vertical axis.
@@
In Fig. 8 are shown $g_{11}(r)$ and $g_{12}(r)$ for $\rho^{\ast}=0.90$ at $T^{\ast}=0.0625$. As seen, both functions oscillate in phase, according to the expectation that a unique underlying disordered equilibrium configuration of the ideal-glass exists, around which the atoms of the two replicas vibrate. Moreover, $g_{12}(r)$ signals as well the local structural change occurring within the ideal-glass, with the presence of the additional peak at about $r/\sigma=1.45$, in between the original first two.
@@
To close this discussion, HMSA/WCA $g(r)$ at $\rho^{\ast}=0.90$ and $T^{\ast}=0.2$ is shown as a dashed line in Fig. 7A. As seen, it compares well with ODS $g(r)$, but in the proximity of the first minimum. This discrepancy will progressively transform, as the temperature is lowered, into a negative “correlation hole”, eventually triggering the failure of the HMSA/WCA numerical integration algorithm. Finally, the study in Ref. 74 offers us the opportunity to positively assess our theoretical predictions for $g(r)$ against molecular dynamics results, at least in one case. This comparison is shown at $\rho^{\ast}=0.95$ in Fig. 7B, for $T^{\ast}=0.233$, i.e. the lowest temperature attained in that paper. Looking at small discrepancies in the figure, we observe that the deformation of the second peak, barely arising in the MD $g(r)$, turns out to be simply delayed to lower temperatures by HMSA. We stress once again that — to the best of our knowledge — no much more simulation structural data are available in the low-temperature regime to compare with, for the monodisperse Lennard-Jones model.